13. Suppose that the demand and supply equations have been estimated and that the demand and supply curves are given by Qd = 14 - 2P & Qs = 2+ 4P; Determine the equilibrium price and quantity
1
Expert's answer
2020-11-06T10:10:22-0500

"\bold {Answers}"

Equilibrium quantity, "Q_{e}=" "10 \space units"


Equilibrium price, "P_{e}=" "\$2"


"\bold {Solutions}"

Market equilibrium is achieved when demand equals supply, that is when

"Qd = Qs"

"=> 14-2P = 2 + 4P"

"=> 14-2 = 4P + 2P"

"12 = 6P"

"\dfrac {6P}{6}=\dfrac {12}{6}"

"\bold {P = \$2}"



When "P = \$2, \space Qd = Qs"

"=> Q =14-2(2)"

"Q = 14-4"


"\bold {Q = 10 \space units}"


"\therefore" the equilibrium price is "\$2" and the equilibrium quantity is "10 \space units"